OpenStudy (anonymous):

A college ended six straight years of tuition increases by raising its tuition from $3000 to $3100. Find the percent increase (round to the nearest tenth of a percent).

OpenStudy (anonymous):

3 %?

OpenStudy (anonymous):

nope 3.3% thanks though

OpenStudy (anonymous):

\[\text{Percent increase}=\left(\frac{\text{Final value}-\text{Initial value}}{\text{Initial value}}\right) \cdot 100\%\]